OpenStudy (anonymous):

Giving MEDALS, please help answer these few questions please! Your friend has $100 when he goes to the fair. He spends $10 to enter the fair and $20 on food. Rides at the fair cost $2.00 per ride. Which function can be used to determine how much money he has left over after x rides? f(x) = -2x + 70 f(x) = 2x + 70 f(x) = -2x - 70 f(x) = -30x + $100

OpenStudy (anonymous):

A State Park charges $5 per car plus $1 per person as an admission fee. The total charged for a car with x people is f(x) = x + 5. How will the graph of this function change if the per car charge is changed to $8 per car? The line will shift vertically down by $3. The line will shift vertically up by $3. The line will shift vertically up by $8. The line will shift vertically down by $8.

OpenStudy (anonymous):

Could someone help me with these two questions, I'm giving medals

OpenStudy (anonymous):

notice me senpai ;-;

OpenStudy (anonymous):

for the second B. the line will shift vertically up by $3.

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Thank you so much @ActionSPORT902100

OpenStudy (anonymous):

Could you possibly help me with another one?

OpenStudy (anonymous):

So, we first take the deduction from the initial $100. Deducting 20 and 10 respectively, we're left with $70. Now, the $2 is going to have a subtraction effect on the remaining $70. Our answer is: x tickets times -2 + the $70 he had before the rides! Or option A.

OpenStudy (anonymous):

no prob:)

OpenStudy (anonymous):

why are my things moving!:(and i'll be happy to:)

OpenStudy (anonymous):

Two sisters decide to take a series of acting lessons. One decides to pay a flat rate of $75, shown as f(x). The other just wants to pay $5 per lesson, shown as g(x). Which function shows the correct combination of these two functions to represent the amount the parent should pay, shown as h(x)? f(x) = 75, g(x) = 5x, h(x) = 5x + 75 f(x) = 75x, g(x) = 5, h(x) = 75x + 5 f(x) = 75, g(x) = 5, h(x) = 5 + 75 f(x) = 75x, g(x) = 5x, h(x) = 75x + 5x

OpenStudy (anonymous):

f(x) = 75, g(x) = 5x, h(x) = 5x + 75 f(x) = 75x, g(x) = 5, h(x) = 75x + 5 f(x) = 75, g(x) = 5, h(x) = 5 + 75 f(x) = 75x, g(x) = 5x, h(x) = 75x + 5x

OpenStudy (anonymous):

A.) F(x)=75,G(x)=5x,H(x)=5x+75

OpenStudy (anonymous):

Thanks again, I have one more question. If you could please help me with?